OpenStudy (anonymous):

WILL REWARD!! Write an equation for the translation of y=5/x that has the asymptotes x=6 and y=7.

hero (hero):

You are given y = 5/x and asked to find asymptotes x = 6 and y = 7 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Hero

hero (hero):

Okay, so the parent function is y = 1/x correct?

OpenStudy (anonymous):

yes

hero (hero):

If you recall, the asymptotes of y = 1/x is x = 0 and y = 0. Do you agree?

OpenStudy (anonymous):

yes i agree

hero (hero):

Okay, so @mrivera3113, do you think multiplying the right side by 5 would change the asymptotes in any way? In other words, are the asymptotes of y = 5/x still x = 0 and y = 0?

hero (hero):

Something else to consider: We can write a formula that represents translations of parente function y = 1/x: \[y = \frac{a}{x \pm b} \pm c\] Where a = vertical stretch or compress b = the value of the x asymptote c = the value of the y asymptote At this point, we already have \[y = \frac{5}{x \pm b} \pm c\] We just have to figure out the values of b and c, which are really already given to us.

hero (hero):

@mrivera3113, I assume you have figured out the correct result for this?